Monday, June 20, 2005

ODOT 156 Steel Fit Up

ODOT sent back an initial response to RFI 029 that we would not be able to
make field cuts of pan-steel.

Sunday, June 19, 2005

ODOT 156 Steel Plates RFI

The engineer has reviewed the request to field trim the steel and has sent
his response to ODOT. Unfortunately, we were not included on the response.
With any luck, we'll be informed of the response on Monday, June 20.

Friday, June 17, 2005

7-Mile Remobilization

Contract Dewatering confirmed that they would have their casings on site on
Tuesday next week and the rig would be there Tue PM or Wed AM. Sunesis will
be mobing an excavator to the site today and anticipated "undoing" the
existing pipes on Monday and Tuesday and straightening up/grading the site
for their rig.

Lburg Pump Station

At 7-Mile, Electrician will be working this weekend. Flygt is set up for
either Wed or Thur next week. City agreed that it will be very difficult to
test larger pumps (as there is insufficient flow or availability of water).
City is supposed to look for electrical connectors today (Friday). If they
can't find them then they stated that we should just order them. These
connectors are not a small dollar item.

Thursday, June 16, 2005

7-Mile - Re-Mobing

We should be remobilizing equipment to the site tomorrow (Friday). Our plan
is to begin dismantling the existing outfall piping (the single 16" pipe
won't handle the new flow) and reinstalling two new 16" next week. Our
Dewatering sub is scheduled to be on site on Monday or Tuesday.

WRB Texas Rail Embedment

Manufacture's recommendation for embedment is 10-15 times the diameter. For
our #5 bar 15 times is 75/8" or 9.375". We'll drill the holes for the Texas
Rail no shorter than 9 inches.

Lburg 06.16.2005

Electricians will be working this weekend. Seagram's road should be back in
service today. Site passed safety-inspection.

7. Mile 06-16-2005

Contract Dewatering called and will now be able to re-mob next week.

Wednesday, June 15, 2005

ODOT 156 06.15.2005

Bridge Steel was set last week. Steel Erector has finished bolting the
structure up. Would like to continue working night shift next Monday
evening through next Thursday evening. LEOs will again be required.

Forming of backwall is progressing -- issues are open.

Field visit is required.

WRB 06.15.2005

Deck is curing. Pour went well. Lessons learned - requested greater rate
-- especially the first hour before the bidwell moves. Ensure the correct
number of people are on site -- not too few, not too many. Get a "real good
feel" for final truck amount needed. Two final orders are too many.

Need to line up asphalter and check on doing base and intermediate, then

Survey markers need installed.

Texas Rail formliners - don't want to modify or tweak forms -- they should
be usable - as ordered. If not, then there is an issue with the supplier.

7-Mile 06.15.2005

Dropped by the site today. Water is just below the lower ranger. We'll
remove all pipes on the southside and regrade for drill rig to access site
from inside.

Received permission from land owner to access his property (if needed --
we'll barter doing some skidsteer work). Spoke to Contract Dewatering -
rigs are tied up, will now try to mobilize by June 27th (week of).

Sluice Gate should ship July 31st.